1 point) It is known that a certain lacrosse goalie will successfully make a save 89 84% of the time. Suppose that the lacrosse goalie attempts to make 13 saves. What is the probability that the lacrosse goalie will make at least 10 saves? Let X be the random variable which denotes the number of saves that are made by the lacrosse goalie. Find the expected value and standard deviation of the random variable E(X) Note: you can find special formulas for the mean and standard deviation of a binomial random variable in our text it you dont want to do this the way we did in Section 7.4
Answer #1

1) X follow binomial distribution with n = 13 , p = 0.8984

P(X >= 10) = 0.9640

2)

E(X) = np = 13 * 0.8984 = 11.6792

sd(S) = sqrt(npq) = sqrt( 13 * 0.8984 * (1 - 0.8984) = 1.08931

Please rate